OpenStudy (anonymous):

Calc 1 question...need to find derivative of 1/sqrt x I know the answer is -(1/2x^3/2) but no clue how to show the work - I suck when radicals and fractions together. I need to use the (f(x+h)-f(x))/h method. Any help would be greatly appreciated =]

OpenStudy (anonymous):

you have to do this by hand right? using \[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\]?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

I've done like 20 problems already but none involved a radical/fraction combo like this

OpenStudy (anonymous):

start out with \[f(x+h)-f(x)=\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

then subtract, just like you would with normal fractions to get \[\frac{\sqrt{x}-\sqrt{x+h}}{\sqrt{x}\sqrt{x+h}}\]

OpenStudy (anonymous):

now rationalize the numerator, multiply top and bottom by \(\sqrt{x}+\sqrt{x+h}\) leave the denominator in factored form

OpenStudy (anonymous):

getting a bit confused reading the text as opposed to seeing it on paper, would you mind if i upload image of what I have down so far on paper? I think I've done what you have said but stuck on last part rationalizing the numerator

OpenStudy (anonymous):

by which i mean write \[\frac{(x-x-h)}{\sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})}\]

OpenStudy (anonymous):

i hope that last step was clear then \(x-x=0\) so numerator is just \(-h\) divide all this by \(h\) and the numerator becomes \(-1\)

OpenStudy (anonymous):

then replace \(h\) by 0 and you are done

OpenStudy (anonymous):

since that picture i multipled top and bottom by sqrt (x+h) +sqrt x

OpenStudy (anonymous):

that is the step before you rationalize the numerator

OpenStudy (anonymous):

ok right, then multiply as you said

OpenStudy (anonymous):

but i would write it as multiplying by \(\sqrt{x}+\sqrt{x+h}\) of course it doesn't matter, but it is easier to see what you will get

OpenStudy (anonymous):

got yah

OpenStudy (anonymous):

okay so the top should be canceled out now with rationalizing...just confused on what the bottom should look like now

OpenStudy (anonymous):

wish there was a way for both people to draw on here so I could see in real time...deciphering the steps is throwing me off a little

OpenStudy (anonymous):

are you having trouble viewing the code? it should be look just like math

OpenStudy (anonymous):

yeah its showing up as \frac{(x-x-h)}{\sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})}\]

OpenStudy (anonymous):

\[\frac{(x-x-h)}{\sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})}\]is what you should have, without the \(h\) in the denominator you get \[\frac{-1}{\sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})}\]

OpenStudy (anonymous):

refresh your browser

OpenStudy (anonymous):

now wonder it was hard to read

OpenStudy (anonymous):

wow there we go lo

OpenStudy (anonymous):

now I can see, let me try now :D

OpenStudy (anonymous):

now it should be clear review the steps, i think they are all there

OpenStudy (anonymous):

just unclear how you got (x-x-h) on the numerator? I thought when you rationalize the numerator it becomes 1?

OpenStudy (anonymous):

\[(\sqrt{x}-\sqrt{x+h})(\sqrt{x}+\sqrt{x+h})=x-(x+h)=-h\]

OpenStudy (anonymous):

then when you divide by \(h\) you get a \(-1\) in the numerator

OpenStudy (anonymous):

ah okay I see now, wow thank you so much for the help...just started calculus 1 last week after just having finished pre-cal this summer so this derivative stuff is killing me :(...always sucked with radicals aswell

OpenStudy (anonymous):

is there a reason though that wolfram shows the answer as\[-(\frac{ 1 }{ 2x ^{3/2} })\]

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

you end up with \[\frac{-1}{\sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})}\] and then let \(h=0\) and get \[\frac{-1}{\sqrt{x}\sqrt{x}(\sqrt{x}+\sqrt{x})}\] \[\frac{-1}{2x\sqrt{x}}=-\frac{1}{2x^{\frac{3}{2}}}\]

OpenStudy (anonymous):

oh! okay

OpenStudy (anonymous):

thank you very much :D